OpenStudy (anonymous):

Group A consists of X students and their total age is 221 and their average age is an integer.When group A is merged with Group B with twice the number of students (the number of students between 30 and 40) average age of B is reduced by 1.What is the original average age of B?

OpenStudy (anonymous):

\[221=13\times 17\] so there are either 13 students and the average age is 17 or there are 17 students and the average is 13. but since group B has twice the number and it is bigger than 30 we know group A has 17 students and average age is 13

OpenStudy (anonymous):

therefore group B has 34 students, total number of students is 51 put total age in group B as x, then the equation to solve is \[\frac{221+x}{51}=\frac{x}{34}-1\]

OpenStudy (anonymous):

get x = 544 and so average os \[\frac{544}{34}=16\]